OpenStudy (anonymous):

How many 5/8 are in 3

OpenStudy (ipwnbunnies):

How many 2's are in 8?

OpenStudy (ipwnbunnies):

Or just ignore my reply.

OpenStudy (texaschic101):

3/(5/8) = 3 * 8/5 = 24/5 = 4 4/5

sammixboo (sammixboo):

Yes it is 4 4/5 like @texaschic101 said.

OpenStudy (texaschic101):

check.. 24/5 * 5/8 = 3 (correct)